OpenStudy (anonymous):

Join (x-27)/(x^(1/3)-3)

OpenStudy (kittiwitti1):

"Join"?

OpenStudy (anonymous):

I don't know what that means, I was hoping someone would. But I googled it and all it took me to was to joining fraction. A website gave me this answer\[x ^{2/3}+3\sqrt[3]{x}+9\] but I don't know how they got there

OpenStudy (kittiwitti1):

hmm I'll type it out in a more readable format first then\[\frac{\left(x-27\right)}{\left(x^{\frac{1}{3}}-3\right)}\]is this right?

OpenStudy (anonymous):

yes

OpenStudy (kittiwitti1):

Okay... do you know how to simplify the exponent 1/3 of the bottom portion?

OpenStudy (anonymous):

i tried multiplying by the bottom conjugate but i got very lost.

OpenStudy (anonymous):

by cubing it

OpenStudy (kittiwitti1):

https://www.mathsisfun.com/algebra/exponent-fractional.html Have you tried this?

OpenStudy (anonymous):

i went on the page and i get the information, I just don't get how to tie it in with the problem

OpenStudy (kittiwitti1):

Well let's convert the x^(1/3) first.

OpenStudy (kittiwitti1):

\[x^{\frac{a}{b}}=\sqrt[b]{x^{a}}\]\[x^{\frac{1}{3}}=\sqrt[3]{x^{1}}\]

OpenStudy (anonymous):

ok i get that part

OpenStudy (kittiwitti1):

Okay so now we have\[\frac{(x-27)}{\sqrt[3]{x}-3}\]

OpenStudy (kittiwitti1):

Hmmm, does this help? http://openstudy.com/study#/updates/4e735b990b8b247045d029a1

OpenStudy (anonymous):

the problem is \[(x-27)\div(\sqrt[3]{x}-3)\] but it simplifies to \[\sqrt[3]{x}^{2}+3\sqrt[3]{x}+9\], i dont understand how to simplify it

OpenStudy (kittiwitti1):

Try the link I posted.

OpenStudy (kittiwitti1):

Does it help?

OpenStudy (anonymous):

yes but i didn't know even without the cubed root with x it qualified as the sum of cube formula

OpenStudy (kittiwitti1):

http://prntscr.com/8g4uw1 Mathematical terminology of "join": to connect with a straight line or curve. Does this help?

OpenStudy (anonymous):

the regular cube formula is\[(x-3)(x ^{2}+3x+9)\]right what i didn't understand is that you can cube root it so it would turn out to \[(x ^{1/3}-3)(x ^{2/3}+x ^{1/3}3+9\]

OpenStudy (kittiwitti1):

@ganeshie8 (Sorry, I'm a tad muddled in the brain here. I may need correcting) From what I can understand they simply substituted the exponent for an x^(1/3)...?

OpenStudy (anonymous):

yes because \[x ^{3}-27\] is a perfect cube and i'm assuming that x-27 is what it look like when its cubed rooted

OpenStudy (kittiwitti1):

Apparently they found a way to use exponent 1/3 to make the process easier.

OpenStudy (kittiwitti1):

\[\frac{x-27}{\sqrt[3]{x}-3}\rightarrow\frac{x^{1/3}-3{x^{1/3}+3x^{1/3}+9}}{\sqrt[3]{x}-3}\]...I think I messed up somewhere.

OpenStudy (anonymous):

you dont distribute it, you would cancel then \[x ^{1/3}-3\] from the top and bottom

OpenStudy (anonymous):

so it would be \[((x ^{1/3}-3)(x ^{2/3}+(3timesx ^{1/3})+9)div(x ^{1/3}-3)\]

OpenStudy (kittiwitti1):

Sorry my mom was cussing me out for being on OS helping people with math ^^;

OpenStudy (kittiwitti1):

It may be worth noting that \[\sqrt[3]{27}=3\]Reason:\[27=3\times3\times3\]I'm stuck here, sorry. Maybe ask @ganeshie8 for help e_e!